OpenStudy (anonymous):

Ben has four more than twice as many CDs as Jake. If they have a total of 31 CDs, how many CDs does Jake have? (1) 9 (3) 14 (2) 13 (4) 22

OpenStudy (anonymous):

I got the answer equation 1) Ben = 4 + 2*Jake equation 2) 31 = Ben + Jake substitute equation 1 into equation 2 31 = (4 + 2*Jake) + Jake 27 = 2*Jake + Jake 27 = (2 + 1)*Jake 27 = 3*Jake 9 = Jake